Calculate ΔSsurr at the indicated temperature for a reaction.

ΔH∘rxn= 126 kJ ; 66 K

Express your answer using two significant figures.
Answer #1

Given:

ΔH = 126 KJ

= 1.26*10^5 J

T = 66 K

use:

ΔS Surr = -ΔH/T

= -(1.26*10^5)/66

= -1.909*10^3 J/K

Answer: -1.9*10^3 J/K